# Vision Transformers for End-to-End Vision-Based Quadrotor Obstacle Avoidance

### v2

@@ I INTRODUCTION
This work’s aim is to study the ViT as an end-to-end framework for quadrotor control, with comparisons to baseline architectures. We choose to train models via behavior cloning in simulation from a privileged obstacle-aware expert, as is common among existing learning-to-dodge works [loquercio2018dronet, Loquercio_2021]. This additionally eliminates potentially confounding factors that may exist in a reinforcement learning framework, such as the exploration/exploitation tradeoff and additional hyperparameter tuning. We use depth sensing data to train and evaluate these models, as across multiple sensor types, robotic platforms, and environments, depth data has shown in prior work to be both ubiquitous and effective at high-level scene understanding as well as navigation tasks. We evaluate models against each other by comparing collision metrics during forward flight through a cluttered scene, across trials with increasing forward velocity. We further present the paths taken by each model through a given simulated scene, the command characteristics of each model (acceleration and energy cost), and compare our end-to-end approach to modular baseline methods. Representative hardware experiments further demonstrate agile avoidance around multiple obstacles and at high speeds (Figure 1).
@@ II RELATED WORK
Vision transformers for control. Some research has used a combination of architectures to surpass the performance of individual components for the object detection task. Swin Transformer [liu2021swin] and a ResNet-based module [he2016resnet] have been combined previously [hendria2023combining], showing that this attention+convolutional network architecture improved object detection performance. We similarly explore combination models between basic backbone architectures (convolution and U-Net, recurrence, and attention) in this work. When looking at downstream tasks such as control from pixels, transformer architectures such as ViT have been compared to convolutional neural networks (CNNs) and shown to perform worse than the CNN architecture due to the weaker inductive bias and need for more training data [tao2022evaluating, li2023a]. ViT has been studied extensively for its use in downstream control tasks as a representation learner in many reinforcement learning (RL) works [hansen2021stabilizing, kalantari2022improving, seo2023masked]. Our method investigates leveraging similar architectures using imitation learning from a privileged expert as opposed to RL.
@@
Quadrotor flight from depth images. When navigating in uncertain environments, learning from depth images offers an input modality for reactive control [Loquercio_2021, zhou2019robust, florence2020integrated]. Various approaches have been taken to learn control outputs from depth image inputs such as mapping then planning [zhou2019robust], using motion primitives [florence2020integrated], and sampling collision-free trajectories [Loquercio_2021]. These methods use depth images to determine some intermediate representation which is then used to control the quadrotor. Compared to these methods, we demonstrate purely end-to-end control using various architectures learning from depth images directly, without the need for mapping, planning, or trajectory generation.
@@ III-B Learning framework
The student model takes as input a single depth image $\mathbf{im}_{depth}\in[0,1]^{60\times 90}$ (similar to [Loquercio_2021]) of size 60$\times$90, quadrotor attitude $\mathbf{q}_{att}=(w,x,y,z)$, and forward velocity $v_{fwd}\in\mathbb{R}$, and predicts a velocity vector $\mathbf{v}_{pred}\in\mathbb{R}^{3}$. As models have access to the forward velocity, we do not include variations of model input such as stacked image sequences; while this input might improve temporal predictions of convolution-only models, we instead present recurrent models that are meant to model sequential data. The $\mathbf{v}_{pred}$ are learned from the privileged expert $\mathbf{v}_{cmd}$ in a supervised behavior cloning fashion, with a standard L2 loss over all timesteps in a trajectory, given student model parameters $\mathbf{\theta}$.
@@ III-C Simulated setup and models
Privileged expert policy and data gathering. To gather supervisory data pairs of $\{(\mathbf{im}_{depth},\mathbf{q}_{att},v_{fwd}),\mathbf{v}_{pred}\}$ for behavior cloning, the privileged expert runs its short-horizon collision avoidance algorithm through randomized obstacle fields. Note that bold variables are multi-dimensional. In total, 588 expert runs yielded 112k depth images, where 27.5k images are sampled from trials where there is at least one collision. Since the expert is reactive and lacks dynamics-feasible planning, collisions are not absent from the data; however, we find that some models (see Section IV-A) outperform this expert at high speeds.
@@
End-to-end student models. We present five student models for a comparison of ViT-based models against other state-of-the-art image processing architectures. Each model takes as input a single depth image, and uses an encoder-like module composed of either convolutions or attention to generate a lower-dimensional middle representation, which is then concatenated with the quadrotor attitude (quaternion) and forward velocity (scalar) before feeding into either LSTM or fully-connected layers. With the exception of the ConvNet model, all models are roughly $\sim$3M parameters (Table I). This model size ensures fast enough computation for in-flight processing in a robotics control loop (30Hz on the hardware platform presented in Section III-D). Certain models such as a fully transformer-based architecture were attempted, but performed poorly and are not presented (unless a fully-connected head was used, as is described below as ViT).
@@
ConvNet+LSTM. A multi-layer LSTM module placed between the convolutional and fully connected layers is included to encourage temporal consistency in velocity predictions.
@@ IV-G Real world experiments
Figure 7: Representative real-world experiments using the Falcon250 (7(a) and Section III-D) evaluating the ConvNet and ViT+LSTM models in the same single, rigid obstacle avoidance task (7(b)). Paths are plotted in 7(c), where lighter shades on the red (ConvNet) and blue (ViT+LSTM) trajectories represent higher lateral velocities. ViT+LSTM consistently avoids earlier than the ConvNet model.
We zero-shot deploy ViT+LSTM in single and multi-obstacle configurations requiring collision avoidance and flight-through-gap behaviors at high speeds up to 7m/s. The model runs inference onboard at 30Hz (CPU) during all trials. Figure 1 shows representative examples of two such trials, with 1(c) showing the depth image input to the model with a red arrow representing the model’s corresponding output velocity command. The quadrotor avoids obstacles in the real world with no model re-training or extensive system fine-tuning; while depth camera images exhibit artifacts not seen in the simulated training data, such as high-frequency patterns, varying background appearance, and propellers in-view (uncropped depth images viewable on project website) our end-to-end learned models performed effectively in many scenes at speeds of 1-7m/s. Figure 7 additionally demonstrates real-world differences in collision avoidance between the ConvNet and ViT+LSTM models.
@@ V CONCLUSION
ACKNOWLEDGMENTS
This work was funded by NSF SLES-2331880 and NSF CAREER-2045834.

### v3

@@ I INTRODUCTION
This work’s aim is to study the ViT as an end-to-end framework for quadrotor control, with comparisons to baseline architectures. We choose to train models via behavior cloning in simulation from a privileged obstacle-aware expert, as is common among existing learning-to-avoid works [loquercio2018dronet, Loquercio_2021]. This additionally eliminates potentially confounding factors that may exist in a reinforcement learning framework, such as the exploration/exploitation tradeoff and additional hyperparameter tuning. We use depth sensing data to train and evaluate these models, as across multiple sensor types, robotic platforms, and environments, depth data has shown in prior work to be both ubiquitous and effective at high-level scene understanding as well as navigation tasks. We evaluate models against each other by comparing collision metrics during forward flight through a cluttered scene, across trials with increasing forward velocity. We further present the paths taken by each model through a given simulated scene, the command characteristics of each model (acceleration and energy cost), and compare our end-to-end approach to modular baseline methods. Representative hardware experiments further demonstrate agile avoidance around multiple obstacles and at high speeds (Figure 1).
@@ II RELATED WORK
Vision transformers for control. Some research has used a combination of architectures to surpass the performance of individual components for the object detection task. Swin Transformer [liu2021swin] and a ResNet-based module [he2016resnet] have been combined previously [hendria2023combining], showing that this attention+convolutional network architecture improved object detection performance. We similarly explore combination models between basic backbone architectures (convolution and U-Net, recurrence, and attention) in this work. When considering downstream tasks such as control from pixels, transformer architectures such as ViT have been compared to convolutional neural networks (CNNs) and shown to perform worse than the CNN architecture due to the weaker inductive bias and need for more training data [tao2022evaluating, li2023a]. ViT has been studied extensively for its use in downstream control tasks as a representation learner in many reinforcement learning (RL) works [hansen2021stabilizing, kalantari2022improving, seo2023masked]. Our method investigates leveraging similar architectures using imitation learning from a privileged expert as opposed to RL.
@@
Quadrotor flight from depth images. When navigating in uncertain environments, depth images offer an input modality for learned, reactive control [Loquercio_2021, zhou2019robust, florence2020integrated] and can even be treated as an intermediate, learnable task for downstream control [bhattacharya2024monocular]. Various approaches have been taken to learn control outputs from depth image inputs such as mapping then planning [zhou2019robust], using motion primitives [florence2020integrated], and sampling collision-free trajectories [Loquercio_2021]. These methods use depth images to determine some intermediate representation which is then used to control the quadrotor. Compared to these methods, we demonstrate purely end-to-end control using various architectures learning from depth images directly, without the need for mapping, planning, or trajectory generation.
@@ III-B Learning framework
The student model takes as input a single depth image $\mathbf{im}_{depth}\in[0,1]^{60\times 90}$ (similar to [Loquercio_2021]) of size 60$\times$90, quadrotor attitude $\mathbf{q}_{att}=(w,x,y,z)$, and forward velocity $v_{fwd}\in\mathbb{R}$, and predicts a velocity vector $\mathbf{v}_{pred}\in\mathbb{R}^{3}$ (see Figure 2). As models have access to the forward velocity, we do not include variations of model input such as stacked image sequences; while this input might improve temporal predictions of convolution-only models, we instead present recurrent models that are meant to model sequential data. The $\mathbf{v}_{pred}$ are learned from privileged expert $\mathbf{v}_{cmd}$ via supervised behavior cloning, with a L2 loss over all timesteps in a trajectory, given student model parameters $\mathbf{\theta}$.
@@ III-C Simulated setup and models
Privileged expert policy and data collection. To gather supervisory data pairs of $\{(\mathbf{im}_{depth},\mathbf{q}_{att},v_{fwd}),\mathbf{v}_{pred}\}$ for behavior cloning, the privileged expert runs a short-horizon collision avoidance algorithm through randomized obstacle fields. Note that bold variables are multidimensional. In total, 588 expert runs yielded 112k depth images, where 27.5k images are sampled from trials where there is at least one collision. Since the expert is reactive and lacks dynamics-feasible planning, collisions are not absent from the data; however, we find that some models (see Section IV-A) outperform this expert at high speeds.
@@
End-to-end student models. We present five student models for a comparison of ViT-based models against other state-of-the-art image processing architectures. Each model takes as input a single depth image, and uses an encoder-like module composed of either convolutions or attention to generate a lower-dimensional middle representation, which is then concatenated with the quadrotor attitude (quaternion) and forward velocity (scalar) before feeding into either LSTM or fully-connected layers. With the exception of the ConvNet model, all models are roughly $\sim$3M parameters (Table I). This model size ensures fast enough computation for in-flight processing in a robotics control loop (30Hz on the hardware platform presented in Section III-D). Certain models such as a fully transformer-based architecture were attempted, but performed poorly and are not presented (unless a fully-connected head was used, as is described below as ViT). Architecture details are online [bhattacharya2024vision].
@@
LSTMnet. A multi-layer LSTM module placed between the convolutional and fully connected layers is included to encourage temporal consistency in velocity predictions.
@@ IV-G Real world experiments
Figure 7: Representative real trials using the Falcon250 (7(a)) evaluating ConvNet and ViT+LSTM in the same single, rigid obstacle avoidance task (7(b)). Paths are plotted in 7(c), where lighter shades on the red (ConvNet) and blue (ViT+LSTM) trajectories represent higher lateral velocities. ViT+LSTM consistently avoids earlier than the ConvNet model.
We zero-shot deploy ViT+LSTM in single and multi-obstacle configurations requiring collision avoidance and flight-through-gap behaviors at high speeds up to 7m/s. The model runs inference onboard at 30Hz (CPU) during all trials. Figure 1 shows representative examples of two such trials, with 1(c) showing the depth image input to the model with a red arrow representing the model’s corresponding output velocity command. The quadrotor avoids obstacles in the real world with no model re-training or extensive system tuning; while depth camera images exhibit artifacts not seen in the simulated training data, such as high-frequency patterns, varying background appearance, and propellers in-view (uncropped depth images viewable on project website), end-to-end learned models performed effectively in many scenes at speeds of 1-7m/s. Figure 7 additionally demonstrates real-world differences in collision avoidance between the ConvNet and ViT+LSTM models.